# Coarse-to-fine Alignment Makes Better Speech-image Retrieval

### v1

@@ I Introduction
With the development of VGS models, the accuracy of speech-image retrieval systems has also significantly improved. This showcases that speech-image retrieval holds great appeal as a standalone application. In FaST-VGS [8], authors employ an innovative training and retrieval approach that combines the dual-encoder and cross-attention architectures, resulting in a single model that achieves both speedy and precise speech-image retrieval capabilities. SpeechCLIP [9] utilizes a speech encoder, initialized with a pre-trained speech self-supervised learning (SSL) model, to align with a frozen CLIP [10] image encoder using paired speech-image data. This alignment of the speech and image embedding spaces enables SpeechCLIP to achieve state-of-the-art performance in speech-image retrieval tasks.
While these methods have proven to be effective, they do have certain limitations. For instance, in FaST-VGS, the use of an object detector as the image encoder may restrict its expressive power. As it is confined to the capabilities of the object detector and its predefined visual vocabulary. SpeechCLIP replaces the object detector with CLIP to extract image features. However, it solely relies on contrastive learning tasks to align speech and image features at a coarse level, which could be challenging to achieve satisfactory alignment. For instance, it may occasionally result in false positives when images and speech exhibit similar semantics but vary in intricate details. Moreover, these approaches may be susceptible to noisy data in the training datasets, which can adversely affect their overall generalization performance.
@@ II-A Preliminaries
Hidden-unit BERT (HuBERT) [13]. HuBERT is a self-supervised learning speech model that utilizes a masked prediction objective, similar to the renowned BERT [14] model. It predicts masked speech frames by considering the surrounding context. It comprises a CNN feature extractor followed by a transformer encoder. It can effectively extract valuable speech representations for various downstream tasks [15].
@@ II-B Architecture
As shown in Figure 1, our framework comprises a speech feature extractor, two unimodal encoders, and a multimodal encoder. We employ a self-supervised learning speech model HuBERT [13] to extract speech features. Inspired by SUPERB [17], we combine the CNN output of HuBERT and the hidden representations from its transformer encoder using learnable weights. This weighted sum of HuBERT’s output forms a sequence of speech features. These speech features, along with the CLS token, are then fed to the transformer speech encoder to extract speech embeddings $S=\{S_{cls},S_{1},...,S_{N}\}$. Additionally, we utilize the visual encoder of BLIP-2 [16] to extract image embeddings $I=\{I_{cls},I_{1},...,I_{N}\}$. A transformer multimodal encoder is used to interact between speech and image embeddings. We utilize $S_{cls}$ and $I_{cls}$ to facilitate speech-image contrastive learning. Subsequently, image embeddings $I$ and speech embeddings $S$ are then fed to the multimodal encoder for interaction through speech-image matching learning. To optimize the learning process, we utilize a large image embedding queue that facilitates the incorporation of numerous negative samples during SIC learning. The queue also enables efficient sampling of hard negative examples for the SIM tasks based on contrastive similarities calculated in SIC tasks. Additionally, in order to improve learning with noisy data, we generate pseudo-targets using the momentum model (a moving-average version of the base model) as additional supervision during training.
@@ II-C Training Objectives
Our model is jointly trained with two main objectives: speech-image contrastive learning on the unimodal encoders and speech-image matching on the multimodal encoder.
Speech-Image Contrastive Learning aims to align the speech and the image features at a coarse level, making it easier for the multimodal encoder to perform cross-modal learning. It learns a similarity function $s=S_{cls}^{T}I_{cls}$, such that parallel speech-image pairs have higher similarity scores compared to non-parallel pairs. Here, $S_{cls}$ and $I_{cls}$ represent the normalized speech and image semantic embeddings, respectively. To effectively increase the diversity of negative examples and enhance the model’s ability to discriminate between positive and negative pairs in the contrastive learning process, we utilize a queue to store a fixed size of image embeddings. It is worth noting that, unlike in [11] where the encoder is updated every iteration, our image encoder remains fixed. By leveraging the queue, we can access a larger pool of diverse image embeddings for each training iteration. For each speech, we calculate the softmax-normalized similarity between the speech and image embeddings as follows:
@@
where $B$ is the batch size. Let $\boldsymbol{y}^{\mathrm{s}2\mathrm{i}}(S)$ and $\boldsymbol{y}^{\mathrm{i}2\mathrm{s}}(I)$ represent the ground-truth one-hot similarity, where negative pairs have a probability of 0, and the positive pair has a probability of 1. The speech-image contrastive loss is defined as the cross-entropy $\mathrm{H}$ between $\boldsymbol{p}$ and $\boldsymbol{y}$ as follows:
@@ III-C Zero-Shot Speech-Image Retrieval
In order to evaluate the generalization ability of our framework, we performed zero-shot retrieval by directly assessing the model trained on SpokenCOCO on the testing sets of Flickr8K. To the best of the author’s knowledge, this is the first time the exploration of the generalization capability of the speech-image retrieval model has been proposed. The result is shown in Table II, where Supervised indicates the model trained on Flickr8k training sets. Surprisingly, the model trained on the SpokenCOCO training sets outperforms the model trained on Flickr8k training sets by a large margin. This demonstrates the excellent generalization ability of our model. The superior performance attributes to the model being trained on a larger SpokenCOCO dataset in comparison to the Flickr8k dataset. In other words, our model demonstrates good scalability. We strongly believe that training on a larger corpus will further enhance its generalization capabilities.

### v2

@@ I Introduction
With the development of VGS models, the accuracy of speech-image retrieval systems has also significantly improved. This showcases that speech-image retrieval holds great appeal as a standalone application. In FaST-VGS [8], authors employ an innovative training and retrieval approach that combines the dual-encoder and cross-attention architectures, resulting in a single model that achieves both speedy and precise speech-image retrieval capabilities. SpeechCLIP [9] leverages a speech encoder, initially pre-trained with a self-supervised learning (SSL) model, and aligns it with a frozen CLIP [10] image encoder using paired speech-image data. This process of aligning speech and image embeddings enables SpeechCLIP to excel in speech-image retrieval tasks, setting new performance benchmarks.
Although these methods have shown effectiveness, they come with certain limitations. For example, in FaST-VGS, using an object detector as the image encoder may limit its expressive power due to the constraints imposed by the detector’s capabilities and its predefined visual vocabulary. SpeechCLIP substitutes the object detector with the CLIP image encoder to extract image features. However, it primarily relies on contrastive learning tasks to align speech and image features at a high level, which can make achieving precise alignment challenging. This approach can sometimes lead to false positives when the images and speech share similar semantics but differ in finer details. Moreover, these approaches may be susceptible to noisy data in the training datasets, which can adversely affect their overall generalization performance.
@@ II-A Preliminaries
Hidden-unit BERT (HuBERT) [13]. HuBERT is a self-supervised learning speech model that uses a masked prediction objective, similar to the widely recognized BERT [14] model. It predicts masked speech frames by taking into account the surrounding context. This design allows HuBERT to efficiently extract valuable speech representations for a range of downstream tasks [15].
@@ II-B Architecture
As shown in Figure 1, our framework comprises a speech feature extractor, two unimodal encoders, and a multimodal encoder. We employ a self-supervised learning speech model HuBERT [13] to extract speech features. Taking inspiration from SUPERB [17], we fuse the CNN output of HuBERT and the hidden representations from its transformer encoder using trainable weights. This weighted aggregation of HuBERT’s outputs generates a sequence of frame speech features. These features, along with the CLS token, are then fed into the transformer speech encoder to obtain speech embeddings $S=\{S_{cls},S_{1},...,S_{N}\}$. Additionally, we utilize the visual encoder of BLIP-2 [16] to extract image embeddings $I=\{I_{cls},I_{1},...,I_{N}\}$. A transformer multimodal encoder is used to interact between speech and image embeddings. We utilize $S_{cls}$ and $I_{cls}$ to facilitate speech-image contrastive learning. Subsequently, image embeddings $I$ and speech embeddings $S$ are then fed to the multimodal encoder for interaction through speech-image matching learning. To optimize the learning process, we utilize a large image embedding queue that facilitates the incorporation of numerous negative samples during SIC learning. The queue also enables efficient sampling of hard negative examples for the SIM tasks based on contrastive similarities calculated in SIC tasks. Additionally, in order to improve learning with noisy data, we generate pseudo-targets using the momentum model (a moving-average version of the base model) as additional supervision during training.
@@ II-C Training Objectives
Our model undergoes joint training with two core objectives: conducting speech-image contrastive learning on the unimodal encoders and performing speech-image matching with the multimodal encoder.
Speech-Image Contrastive Learning is designed to align the speech and the image features at a coarse level, thereby easing the task of the multimodal encoder in cross-modal learning. It optimizes a similarity function $s=S_{cls}^{T}I_{cls}$, such that matching speech-image pairs receive higher similarity scores than mismatched pairs. Here, $S_{cls}$ and $I_{cls}$ refer to the normalized semantic embeddings of the speech and image, respectively. To effectively increase the diversity of negative examples and enhance the model’s ability to discriminate between positive and negative pairs in the contrastive learning process, we utilize a queue to store a fixed size of image embeddings. It is worth noting that, unlike in [11] where the encoder is updated every iteration, our image encoder remains fixed. By leveraging the queue, we can access a larger pool of diverse image embeddings for each training iteration. For each speech, we calculate the softmax-normalized similarity between the speech and image embeddings as follows:
@@
where $B$ is the batch size. Let $\boldsymbol{y}^{\mathrm{s}2\mathrm{i}}(S)$ and $\boldsymbol{y}^{\mathrm{i}2\mathrm{s}}(I)$ represent the ground-truth one-hot similarity, assigning a probability of 0 to negative pairs and 1 to positive pairs. The speech-image contrastive loss is given by the cross-entropy $\mathrm{H}$ between $\boldsymbol{p}$ and $\boldsymbol{y}$, as follows:
@@ III-C Zero-Shot Speech-Image Retrieval
The generalization capability of a model is a crucial metric. Therefore, in this section, we evaluate the model’s generalization by performing zero-shot retrieval. Specifically, we directly assess the model trained on SpokenCOCO using the testing sets of Flickr8K. The results are presented in Table II, where ”Supervised” refers to the model trained on the Flickr8k training sets. Surprisingly, the model trained on the SpokenCOCO training sets achieves significantly better performance than the one trained on the Flickr8k training sets. This underscores the outstanding generalization capability of our model. This superior performance can be attributed to the larger size of the SpokenCOCO training sets compared to the Flickr8k training sets. We are confident that training on an even larger corpus will further improve its generalization capabilities.